# CROSS-GAiT: Cross-Attention-Based Multimodal Representation Fusion for Parametric Gait Adaptation in Complex Terrains

### v2

@@ V-A Implementation
Table I: Comparison of our proposed method (CROSS-GAiT) against STERLING[6] to evaluate the classification performance of learned representations using a Support Vector Classifier.
V-B Dataset Collection
@@ V-D Testing Scenarios
We evaluate our method’s performance in real-world outdoor test scenarios featuring a combination of different terrains, including terrains out of the training dataset (e.g., mulch, irregular rocks, etc). Each scenario (See Fig. 1 and Fig. 5)is tested with at least 10 trials. A trial will be considered successful only if it navigates through the expected terrain.
Scenario 1 - Concrete, mulch, pebbles, and dense vegetation. (See Fig. 1).
@@
Table II: Improved navigation performance of CROSS-GAiT, compared to other methods. We observe considerable improvement on all three metrics. $\dagger$ denotes the best value when considering only successful trials.
V-E Analysis and Discussion
CROSS-GAiT consistently outperforms all other methods in terms of Success Rate and Time to Reach Goal across all four scenarios. This is due to its ability to dynamically switch between gaits, adapting smoothly across a spectrum of gait parameters rather than being restricted to a fixed set.
@@ V-E Analysis and Discussion
In Scenario 2, both AMCO and ProNav switch to predefined gaits that increase both step height and hip splay while traversing the rocks. In contrast, CROSS-GAiT increases hip splay without significantly raising step height, resulting in lower RMS IMU energy density (less vibration) and lower joint effort. This leads to and faster goal completion due to improved stability and fewer instabilities. GA-Nav avoids the rocky regions due to its fixed semantic constmap. This is not the expected behaviour for a direct comparison in our context.
In Scenario 3, AMCO and ProNav switch to predefined gaits with higher steps and a wider stance, leading to higher joint efforts. GA-Nav doesn’t adjust its gait, causing its legs to occasionally get stuck in the sand. In contrast, CROSS-GAiT adapts dynamically, selecting gait parameters that balance between sand and rocks, demonstrating strong generalization. As a result, CROSS-GAiT outperforms all other methods in terms of all defined metrics. GA-Nav gets stuck before unable to navigate through the grass.
@@
Comparison on Embedding space: In Table I, we compare the terrain classification performance of the embeddings generated by CROSS-GAiT and STERLING [6] using a support vector classifier. CROSS-GAiT achieves higher accuracy, precision, recall, and F1-score. By combining visual and time-series data through cross-attention, CROSS-GAiT generates richer latent representations, enhancing the classifier’s ability to distinguish between diverse terrains.
@@ VI Conclusions, Limitations, and Future Work
In this paper, we present a novel framework to enhance robot locomotion across diverse terrains using multimodal sensory inputs—IMU readings, joint forces, and visual inputs. Our approach leverages cross-attention to fuse these inputs, generating a comprehensive terrain representation that allows dynamic adjustment of gait parameters like step height and hip splay. This enables the robot to navigate complex environments with improved stability, energy efficiency, and reduced time to reach the goal.
The key contributions of our method include the fusion of multimodal inputs through cross-attention, creating a richer terrain understanding for more informed gait adjustments. Additionally, we employ a regressor that allows continuous and adaptive adjustment of gait parameters, enabling generalization to unseen terrains, unlike predefined gait strategies. While the method performs well, optimizing the latent representations for all terrain types remains a challenge. Future work will focus on refining these representations, incorporating additional sensory data, using offline reinforcement learning techniques to learn optimal gaits without the need for labeled data, further enhancing the system’s ability to adapt to diverse terrains.
@@ A-B2 Maximum Change Per Iteration
A-C Data Collection Processs (extending Section V-B)

### v3

@@ V-B Dataset Collection
Table I: Comparison of our proposed method, CROSS-GAiT against STERLING [6], along with ablation studies removing either vision or proprioceptive sensor modalities and an ablation replacing cross attention with MLP-based fusion, to evaluate the classification performance of learned representations using a Support Vector Classifier.
@@ V-D Testing Scenarios
We evaluate our method’s performance in real-world outdoor test scenarios featuring a combination of different terrains, including terrains out of the training dataset (e.g., mulch, irregular rocks, etc). Each scenario (see Fig. 1 and Fig. 5) is tested with at least 10 trials. A trial is considered successful only if it navigates through the expected terrain.
Scenario 1 - Concrete, mulch, pebbles, and dense vegetation. (as in Fig. 1).
@@
V-E Analysis and Discussion
CROSS-GAiT consistently outperforms all other methods in Success Rate and Time to Reach Goal across all four scenarios. This superior performance stems from its ability to dynamically adjust gait parameters in real-time, enabling smooth and continuous adaptation across a spectrum of gaits rather than being constrained to a predefined set.
@@ V-E Analysis and Discussion
Table II: Improved navigation performance of CROSS-GAiT, compared to other methods. We observe considerable improvement in all the four metrics. $\dagger$ denotes the best value when considering only successful trials.
@@
In Scenario 2, both AMCO and ProNav switch to predefined gaits that increase both step height and hip splay while traversing the rocks. In contrast, CROSS-GAiT increases hip splay without significantly raising step height, resulting in lower RMS IMU energy density (less vibration) and lower joint effort. This leads to and faster goal completion due to improved stability and fewer instabilities. GA-Nav avoids the rocky regions due to its fixed semantic cost map. This is not the expected behavior for a direct comparison in our context.
In Scenario 3, AMCO and ProNav switch to predefined gaits with higher steps and a wider stance, leading to higher joint efforts. GA-Nav doesn’t adjust its gait, causing its legs to occasionally get stuck in the sand. In contrast, CROSS-GAiT adapts dynamically, selecting gait parameters that balance between sand and rocks, demonstrating strong generalization. As a result, CROSS-GAiT outperforms all other methods in terms of all defined metrics. GA-Nav gets stuck before being able to navigate through the grass.
@@
Comparison on Embedding space: In Table I, we compare the terrain classification performance of the embeddings generated by CROSS-GAiT and STERLING [6]. CROSS-GAiT achieves higher accuracy, precision, recall, and F1-score. By combining visual and time-series data through cross-attention, CROSS-GAiT generates richer latent representations, enhancing the classifier’s ability to distinguish between diverse terrains.
Ablation on Cross-attention: We replace our cross-attention-based fusion with a concatenation followed by an MLP layer, following the approach in [6], as shown in Table I, to assess the benefits of cross-attention. Our results show that cross-attention improves terrain classification accuracy, precision, recall, and F1-score compared to concatenation-based fusion. This improvement stems from cross-attention’s ability to dynamically align and weight features across modalities, enabling richer and more context-aware representations, whereas MLP-based concatenation applies static transformations, limiting its adaptability to multimodal variations.
Real-time Inference: Our lightweight model achieves an inference rate around 60 Hz on a mobile computing platform equipped with an Intel i7 CPU and NVIDIA RTX 2060 GPU, demonstrating its efficiency for real-time continuous gait adaptation. This validates its feasibility for real-world deployment on mobile robots, ensuring responsiveness and adaptability in dynamic environments.
@@ VI Conclusions, Limitations, and Future Work
In this paper, we present a novel framework to enhance robot locomotion across diverse terrains using multimodal sensory inputs: IMU readings, joint forces, and visual inputs. Our approach leverages cross-attention to fuse these inputs, generating a comprehensive terrain representation that allows dynamic adjustment of gait parameters like step height and hip splay. This enables the robot to navigate complex environments with improved stability, energy efficiency, and reduced time to reach the goal.
The key contributions of our method include the fusion of multimodal inputs through cross-attention, creating a richer terrain understanding for more informed gait adjustments. Additionally, we employ a regressor that allows continuous and adaptive adjustment of gait parameters, enabling generalization to unseen terrains, unlike predefined gait strategies. While the method performs well, optimizing the latent representations for all terrain types remains a challenge. Future work will focus on refining these representations, incorporating additional sensory data, and using offline reinforcement learning techniques to learn optimal gaits without the need for labeled data, further enhancing the system’s ability to adapt to diverse terrains.
@@ A-B2 Maximum Change Per Iteration
A-C Data Collection Process (extending Section V-B)